Four consecutive numbers have a sum of 194. Create an equation and use the equation to solve for the numbers.(47, 48, 49, 50)
Found 2 solutions by richwmiller, Alan3354:
Answer by richwmiller(17219)

x+x+1+x+2+x+3=194
4x+6=194
4x=188
x=47
x+1=48
x+2=49
x+3=50

Answer by Alan3354(69443)

194/4 = 48.5, center
--> 47, 48, 49, 50
Don't make a simple problem into a dissertation.